OpenStudy (anonymous):

is it possible for a rectangle with the perimeter of 52 cm to have an area of 148.75 cm?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Just kidding. Let's see...

OpenStudy (anonymous):

Perimeter = 2A+2B Area = A*B 52=2A+2B 148.75=A*B 52-2A=2B (52-2A) / 2 = B 148.75 = A * ((52-2A)/2) \[148.75 = A (\frac{ 52-2A }{ 2 })\] so......

OpenStudy (anonymous):

\[\frac{ 148.75 \times 2 }{ A } = 52 -2A\]

OpenStudy (anonymous):

(we can do that because it's a rectangle so A it's not zero)

OpenStudy (anonymous):

Gosh I got lost. Not sure :(

OpenStudy (anonymous):

im at a COMPLETE loss lol but thanks

OpenStudy (lgbasallote):

anything is possible actually

OpenStudy (anonymous):

so what exactly could i put as the answer then?

OpenStudy (lgbasallote):

you continue what hookdump was doing \[148.75 = A(\frac{52 - 2A}{2})\] \[\implies 148.75 = A(26 - A)\] \[\implies 148.75 = 26A - A^2\] change to quadratic form \[\implies A^2 - 26A + 148.75 = 0\] use quadratic formula to solve for A

OpenStudy (anonymous):

ok........ whats your answer? this looks wrong (my answer)